Shift report 20010823 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Aug 23 2001 - 08:24:50 EDT

  • Next message: Eun-Joo Kim: "Raw data file (run4347)"

    ------------------------------------------------------------
    Date: Thu Aug 23 08:24:50 2001, Shift: 00:00-08:00
    Supervisor: J.I.Jordre, with: T.Lussand
    Log book#8 pages 5 - 6
    
    Shift Summary:
    ------------------------------------------------------------
    1. Had a long store from 01:45 to 07:45. BRAHMS however had less than half the ZDC rate of the other experiments. Called MCR and got answer that they would either try to steer or dump the beam. No beam was dumped and only a small increase in the rate occured. A little bit later MCR called and asked about our rates. They could not see our rates on their computer system, only on the TV screens. No wonder they didn't manange to steer. 
    2. Run 4344 (started in last shift) with 35.2k trigger 1 events
       Run 4347 with 219.8k trigger 1 events.
    3. When Dana arrived he found that both ZDC low and high gain pulses came too early in the gate. He moved the pulses 20ns into the gate.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Aug 23 2001 - 08:25:47 EDT